### v1

# A Survey on Unlearning in Large Language Models

Abstract.
The advancement of Large Language Models (LLMs) has revolutionized natural language processing, yet their training on massive corpora poses significant risks, including the memorization of sensitive personal data, copyrighted material, and knowledge that could facilitate malicious activities. To mitigate these issues and align with legal and ethical standards such as the “right to be forgotten”, machine unlearning has emerged as a critical technique to selectively erase specific knowledge from LLMs without compromising their overall performance. This survey provides a systematic review of over 180 papers on LLM unlearning published since 2021, focusing exclusively on large-scale generative models. Distinct from prior surveys, we introduce novel taxonomies for both unlearning methods and evaluations. We clearly categorize methods into training-time, post-training, and inference-time based on the training stage at which unlearning is applied. For evaluations, we not only systematically compile existing datasets and metrics but also critically analyze their advantages, disadvantages, and applicability, providing practical guidance to the research community. In addition, we discuss key challenges and promising future research directions. Our comprehensive overview aims to inform and guide the ongoing development of secure and reliable LLMs.
Machine Unlearning, Large Language Models
copyright: noneccs: Information systems Language modelsccs: Security and privacy
1. Introduction
Large Language Models (LLMs) have significantly transformed research paradigms in natural language processing while enabling a diverse array of practical applications. These capabilities arise from training on extensive textual corpora, which allows the models to internalize and encode substantial knowledge within their parameters. However, this capacity also introduces critical risks. For instance, personally identifiable information memorized during training can be extracted through privacy attacks, raising concerns under data protection regulations such as the “right to be forgotten” (Voigt and Von dem Bussche, 2017; Regulation, 2016). Similarly, unauthorized use of copyrighted materials in training data can expose model providers to legal challenges (Wei et al., 2025a). Moreover, LLMs can internalize knowledge that facilitates malicious activities (Li et al., 2024; Lang et al., 2025), and jailbreak attacks can elicit the generation of harmful or illegal content. In light of these concerns, selectively erasing specific knowledge from LLMs has emerged as a necessary step toward enhancing their security, reliability, and regulatory compliance.
One potential solution is to retrain LLMs from scratch after removing problematic data. However, this approach is computationally expensive and impractical for large-scale models. Machine unlearning (Cao and Yang, 2015) offers a more efficient alternative, which aims to develop algorithms to selectively remove the influence of specific training data while preserving the overall performance of the model on retained data. In the context of LLMs, the distinctive autoregressive next-token prediction mechanism (Yao et al., 2024a) has motivated extensive research into unlearning methods specifically designed for these models. This survey narrows its focus to address unlearning techniques tailored for large-scale generative language models, which are predominantly used for generative tasks rather than classification. [^1] Some early LLM unlearning works also considered classification tasks in natural language processing (Pawelczyk et al., 2024; Bhaila et al., 2025; Chen and Yang, 2023), but they are not the focus of this survey. By systematically reviewing more than 180 papers published since 2021222Some articles were retrieved from public repositories such as https://github.com/chrisliu298/awesome-llm-unlearning., this survey aims to provide a comprehensive overview of the definition, methods, evaluations, challenges and future directions in LLM unlearning.
Several existing surveys touch upon LLM unlearning, some of which adopt a broader scope or concentrate on specialized aspects (Cooper et al., 2024; Zhou et al., 2024; Liu et al., 2024c; Xu, 2024; Barez et al., 2025; Zhang et al., 2025b; Si et al., 2023; Qu et al., 2025). Compared to surveys that also focus specifically on LLM unlearning (Geng et al., 2025a; Blanco-Justicia et al., 2025), this work offers a more systematic and comprehensive perspective, with several distinctive contributions in the following paragraph. A detailed comparison is summarized in Table 1.
(1) A novel taxonomy of unlearning methods. We categorize unlearning approaches based on the training stage at which they are applied: training time, post-training, and inference time. This taxonomy offers a clearer organizational structure compared to alternative classifications based on parameter selection, since some full-parameter methods can also be applied to part of the parameters or by incorporating LoRA adapters to apply to extra parameters. (2) Multidimensional analysis of evaluations. Instead of merely enumerating existing datasets and metrics, we provide a multidimensional analysis for both datasets and metrics. For datasets, through a comparison from the perspectives of task format, content, and experimental paradigms, we evaluate the characteristics of 18 existing benchmarks, offering actionable guidance for researchers. For metrics, from the goal of LLM unlearning, we analyze knowledge memorization metrics and their applicability, along with commonly used metrics for model utility, robustness, and efficiency. (3) Discussion of Challenges and Future Directions. We provide an in-depth discussion of current challenges in LLM unlearning and outline prospective research directions. These contributions aim to accelerate progress in the emerging field of LLM unlearning, ultimately contributing to safer and more responsible AI systems.
      Current Survey Feature Comparison
      Cooper et al. (2024); Zhou et al. (2024); Liu et al. (2024c) Generative AI Limited attention to generative LLMs.
      Xu (2024) Traditional and large-scale models
      Barez et al. (2025) Open problems Specific aspects of LLM unlearning.
      Zhang et al. (2025b) Legal perspectives
      Si et al. (2023); Qu et al. (2025) Classification-oriented settings
      Geng et al. (2025a); Blanco-Justicia et al. (2025) LLM unlearning (1) Classify by parameters modified.
(2) Enumerate datasets and metrics.
Table 1. A comparison of our work with existing surveys on LLM unlearning. Most surveys focus on a larger or smaller scope than ours. Two surveys similar to ours focusing on LLM Unlearning adopt different classification criteria and lack detailed analysis on datasets and metrics.
2. Backgrounds
2.1. Machine Unlearning in LLMs
Within the standard framework of machine unlearning, we consider a dataset $\mathcal{D}$ and an original model $\mathcal{M}$, parameterized by $\theta$, trained on $\mathcal{D}$. The subset of training data targeted for removal is denoted as the unlearn set $\mathcal{D}_{u}\subset\mathcal{D}$, while the remainder constitutes the retain set $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{u}$. The objective of machine unlearning is to design an algorithm $\mathcal{U}$ that takes as input the original model $\mathcal{M}$ and the relevant data, and outputs an unlearned model $\mathcal{M}_{u}$. This model $\mathcal{M}_{u}$ is intended to approximate the behavior of a retrained model $\mathcal{M}_{r}$, which is trained exclusively on the retain set $\mathcal{D}_{r}$. An illustration of the unlearning process is provided in Figure 1. In the context of LLM unlearning, we provide specific explanations from two aspects: (1) different types of unlearning request and (2) the goal of unlearning.
Refer to captionOriginal Model $\mathcal{M}$Refer to caption?Unlearned Model $\mathcal{M}_{u}$
Unlearning
algorithm $\mathcal{U}$
Refer to captionRetrained Model $\mathcal{M}_{r}$
Compare
Figure 1. Illustration of an unlearning process. The box below the model represents the composition of the corresponding training set. The unlearn set $\mathcal{D}_{u}$ is represented by the shadow square and the retain set $\mathcal{D}_{r}$ is represented by the white square. An unlearning algorithm is applying on the initial target model to obtain the unlearned model $\mathcal{M}_{u}$. And the unlearned model is expected to approximate the retrained model $\mathcal{M}_{r}$.
2.1.1. Type of Unlearning Request
Sample-level Privacy
Anallise Ivory was born on November 8, 1990, and her Social Security Number is 900-55-1236. (Ramakrishna et al., 2025)
Copyright
“There’s more in the frying pan,” said Aunt Petunia, turning eyes on her massive son. (Shi et al., 2025)
Safety
[…]
Entity-level
Concrete
Entity: Stephen King
Samples: Stephen King is a world-renowned American author of horror, suspense, supernatural fiction, … (Jin et al., 2024)
Abstract
Entity: Brute Force
Samples: Adversaries may use brute force techniques to gain access to … (Lang et al., 2025)
Figure 2. Examples of different requests. We extract some fragments from the unlearn set of the corresponding work. At an entity level, in addition to the entity for unlearning, we also show generated samples of these entity, giving an illustration of converting entity-level unlearning to sample-level unlearning.
The predominant form of unlearning request operates at the sample level, requiring models to forget specific text sequences that contain sensitive information, thereby mitigating privacy (Tian et al., 2024; Ramakrishna et al., 2025), copyright (Eldan and Russinovich, 2024; Shi et al., 2025), or safety risks (Li et al., 2024). Examples of different samples are shown in Figure 2. These sequences may consist of free-form text or structured question–answer pairs, as outlined in Table 4.
Beyond isolated samples, a growing number of work addresses entity-level unlearning, which aims at removing all knowledge associated with a particular entity. Entities may be concrete (e.g., individuals, books) (Jin et al., 2024; Choi et al., 2025) or abstract (e.g., biases, capabilities) (Lang et al., 2025), as depicted in Figure 2. Usually, this task is reduced to sample-level unlearning by constructing a corresponding unlearn set with samples related to the target entity. Compared to sample-level unlearning, it requires not only erasing memorized content but also managing inter-entity correlations, rendering it significantly more challenging.
2.1.2. Goals of Unlearning
In traditional machine unlearning, the principal objective of the unlearned model $\mathcal{M}_{u}$ is to behave indistinguishably from the retrained model $\mathcal{M}_{r}$. Consequently, many evaluation approaches rely on comparisons with the retrained model. However, for LLMs, complete retraining is generally infeasible due to the scale of the training data and the inaccessibility of proprietary datasets to external auditors.
Thus, the retrained model cannot serve as a direct reference in the LLM setting. However, by extrapolating from the principles underlying $\mathcal{M}_{r}$, we can identify the core objectives for unlearning:
Goal of LLM unlearning: the unlearned model should no longer memorize content from the unlearn set while preserving all other content.
Meanwhile, we expect the unlearning algorithm to achieve the above objectives with minimal computational and time overhead. Guided by these goals, numerous studies have proposed corresponding evaluation metrics, which we examine in detail in Section 4.
2.2. Related Topics
Several related research areas exhibit conceptual or methodological overlaps with LLM unlearning, offering valuable insights and transferable techniques. However, their core objectives and problem formulations differ from the LLM unlearning paradigm. Hence, we briefly introduce these adjacent fields to clarify correlations and distinctions in this section, while detailed discussions of these topics fall beyond this survey’s scope.
2.2.1. Memorization and Data Extraction
As formalized by the goal of unlearning in Section 2.1.2, the conceptualization of memorization directly shapes the objectives of unlearning, while the methodology for memorization detection provides an essential diagnostic tool for evaluating unlearning efficacy. There exist multiple definitions of LLM memorization, such as formulations based on counterfactual memorization (Zhang et al., 2023a) and tuple completion (Meng et al., 2022). Among these, extractable memorization (Carlini et al., 2021) is the most prevalent, conceptualizing memorization as content that the model can reproduce under specific prompting conditions. This definition originally involved identifying a precise input prefix to induce the model to output the memorized content, and has evolved into a diverse class of data extraction attacks, employing various input strategies and detection mechanisms (Shin et al., 2020; Zou et al., 2023). Consequently, data extraction attacks serve a dual role: they constitute a critical tool for evaluating unlearning, particularly for assessing the knowledge memorization, while unlearning itself functions as a defensive measure to purge hazardous knowledge and thereby mitigate the risks posed by malicious data extraction attempts.
2.2.2. Knowledge Updating
Knowledge editing and updating are essential for maintaining the long-term efficacy of large language models (LLMs), as they enable the correction of inaccuracies and the integration of new knowledge without requiring full model retraining. LLM unlearning can be viewed as a promising strategy within this domain, with research advancing in two main directions: some studies develop robust, conflict-free parameter update algorithms to facilitate reliable knowledge modification (Jung et al., 2025; Ni et al., 2024; Sun et al., 2024), while others apply unlearning techniques to domain-specific contexts (Yang and Lo, 2024; Fore et al., 2024). Another widely adopted paradigm is model editing, which focuses on local, targeted modifications to specific factual knowledge while preserving the model’s general capabilities and avoiding catastrophic forgetting. A key distinction between model editing and unlearning lies in their objectives: model editing operates with a predefined target knowledge state, whereas unlearning aims to remove or suppress information without necessarily replacing it. Nevertheless, mechanistic insights from model editing techniques, such as knowledge neurons and locate-then-edit approaches (Meng et al., 2022), can inform the design of more precise and interpretable unlearning methods.
2.2.3. Alignment
Alignment seeks to ensure that LLMs behave in accordance with human values and intentions. This objective is dual in nature, involving the guide of models toward generating helpful responses (positive) and preventing them from producing undesirable outputs (negative). A widely adopted approach for positive guidance is reinforcement learning from human feedback (RLHF), which steers models toward desirable behaviors through iterative reward-based optimization (Bai et al., 2022; Ouyang et al., 2022). Complementarily, LLM unlearning has emerged as a critical technique for negative alignment, systematically removing undesirable knowledge or capabilities from models. For example, it has been applied to mitigate social biases (Yu et al., 2023; Dige et al., 2024), eliminate unauthorized content to protect copyright (Wei et al., 2025a), and reduce the risk of leaking sensitive information (Rashid et al., 2025; Geng et al., 2025b). Together, these methods form a cohesive alignment framework that addresses both the promotion of beneficial behaviors and the suppression of harmful ones.
3. Existing unlearning methods
Original model,
data, etc.
Output Refer to caption Trained Model TrainingInference SISA-based 3.1 Training-time Input Output Token Embedding Token Logit Instruction Sample Decode Offset 3.3 Inference-time Fine-tuning Others Objectives Parameters Text Distribution Activation Multiple objective Full Partial Extra Data-based Data-free Arithmetic operations SAE-based PEFT Subspace 3.2 Post-training
Figure 3. Framework of unlearning methods. In typical LLM usage scenarios, a model is first trained on specific datasets, and then is used for inference to generate outputs. The unlearning method can be applied to the training process, the trained model, or the inference stage, corresponding to training-time unlearning (Section 3.1, post-training unlearning (Section 3.2) and inference-time unlearning (Section 3.3).
In typical LLM usage scenarios, a model is first trained on specific datasets from draft or a pretrained base model, and then is used for inference to generate output in some tasks. As illustrated in Figure 3, the unlearning method can be applied to the training process, the trained model, or the inference stage, corresponding to training-time unlearning (Section 3.1), post-training unlearning (Section 3.2) and inference-time unlearning (Section 3.3). In short, Training-Time Unlearning requires adjusting the training process to facilitate unlearning, which is mainly based on SISA training paradigms. Post-Training Unlearning involves altering the trained model, mainly through fine-tuning towards multiple objectives on selected parameters. Inference-time Unlearning aims to achieve unlearning via input or output adjustments, rather than modifying the model parameters.
3.1. Training-Time Unlearning
As the pretraining of LLMs typically involves complex procedures and massive datasets, existing training-time unlearning methods primarily focus on the further training phase of a pretrained base model. These approaches address a setting in which a general base model $\mathcal{M}_{b}$ is adapted for specific downstream tasks, during which it may have memorized sensitive information and thus requires unlearning.
As noted in Section 2.1, the ideal outcome of unlearning is to obtain a retrained model $\mathcal{M}_{r}$. However, full retraining starting from $\mathcal{M}_{b}$ is computationally prohibitive and impractical in real-world scenarios. To alleviate this burden, training-time unlearning techniques, exemplified by SISA (Bourtoule et al., 2021), introduce novel training frameworks that initiate retraining from intermediate states. These methods partition the dataset into multiple subsets and store corresponding checkpoints trained on different subsets. By constraining the influence of data points, retraining can start from specific checkpoints unaffected by the unlearn set, thereby accelerating the unlearning process.
Given the considerable storage and computational overhead associated with maintaining numerous model copies, Bannihatti Kumar et al. (2023) and Chowdhury et al. (2025) integrate supplementary trainable components, applying the SISA principle to fine-tune and preserve only the newly introduced parameters. This strategy substantially reduces the number of parameters requiring updates. Beyond efficiency and performance considerations, Kadhe et al. (2023) examine fairness concerns in SISA-based frameworks and propose FairSISA, which integrates three post-processing bias mitigation techniques.
In summary, training-time unlearning ensures, from a mechanistic standpoint, that the model does not encounter data from the unlearn, thereby providing verifiable guarantees. However, this approach is inapplicable to models that have already been fully trained, which significantly constrains its practical applicability.
3.2. Post-Training Unlearning
The main approach to unlearning is to modify the parameters of trained LLMs, which is commonly referred to as the “post-training” phase. This leads to two crucial questions: (1) How to modify the parameters? (2) Which parameters should be selected for modifying? For question (1), the modification of parameters in most methods is an optimization problem, so we will introduce their objective design in Section 3.2.1. While some methods adopt alternative strategies, we include important ones in Section 3.2.2. For question (2), the modified parameters can be all parameters of the model, part of the parameters, or newly introduced parameters. We will discuss how to select part of the parameters in Section 3.2.3 and how to incorporate new parameters in Section 3.2.4. Note that parameter strategies can be freely combined with the objectives in Section 3.2.1.
3.2.1. Objective Optimization
The core of an unlearning mechanism is the design of its objective for optimization, which dictates how the model’s parameters are adjusted to forget specific knowledge while minimizing the impact on general utility. Different from the unified next-token prediction loss during learning, the design of unlearning objective is quite more diverse. Based on the primary target of the designated objective, we can classify existing methods into three major categories: Text-based, Distribution-based, and Activation-based. Figure 4 compares the general pipeline of different objective categories.
Target
LLM
Output textLikelihood-basedUnlearn set / substitute text(a) Text-based
Target
LLM
Output logitsRef. distributionDivergence-basedUnlearn set(b) Distribution-basedTarget LLMHidden activationRef. activationDistance-basedUnlearn set(c) Activation-based
Figure 4. Objective designs of unlearning methods. The color coding is as follows: blue for text, red for tensors/vectors, orange for loss functions. Text-based and distribution-based methods compute a loss function at the output layer by comparing it to a reference (ref.), in textual and distributional level, respectively. Activation-based methods compute the loss using activations from the hidden layers against a reference.
Text-based: These kind of objectives are most intuitive, which aim at minimize or maximize the predicted likelihood of certain text. A representative baseline is Gradient Ascent (Jang et al., 2023a; Yao et al., 2024b). It reduces the prediction probabilities of forget set samples by directly negating their cross-entropy next-token prediction loss. NPO (Zhang et al., 2024b) employs preference-based objective, introduces a reference model to constrain parameter changes. Another approach (Maini et al., 2024; Mekala et al., 2025) focuses on improve likelihood of “substitute responses”, which reply to a query from the forget set appropriately, without disclosing targeted knowledge. Although simple and effective, these baselines often spill unlearning effect beyond forget set itself, which impairs the model’s overall performance. To alleviate this, WGA (Wang et al., 2024) and FPGA (Feng et al., 2024) introduce methods to apply different weights to various token positions within the sequences of forget set. More methods put attention on generate or select data. Some utilize a external LLM to generate substitute responses relevant to the query in forget set  (Mekala et al., 2025; Sinha et al., 2025; Xu et al., 2025). Patil et al. (2025) and Chang and Lee (2025) proposed methods for selecting core subsets from the original unlearning corpus.
Distribution-based: Text-based targets must provide labels from limited vocabulary, which restricts the optimization space of the model during unlearning. In order to achieve more fine-grained unlearning, some methods aim to make the model’s output distribution converge to a reference distribution that aligns with the unlearning goals. The key of these objectives lies in how to construct reference distribution. ME (Yuan et al., 2025b) uses a uniform distribution over the entire vocabulary. For other methods, reference distribution can establish either by modifying data or manipulating logits. WHP (Eldan and Russinovich, 2024) substitutes the unlearning target with unrelated entities to get general knowledge distribution that does not contain target information. WPU (Liu et al., 2025a) improves upon WHP by incorporating diverse substitute entities, performing entity name restoration, and augmenting input prompts. On the other hand, RKLD (Wang et al., 2025a) taking the difference between the logits of a model finetuned on the unlearning set and that of original model as reference. Similar logits-aware approach is also adopted by Obliviate (Russinovich and Salem, 2025) and PerMU (Wang et al., 2025b). Distribution-based objectives generally shorten distance to target distribution via minimizing divergence. The most popular choice is KL divergence, but there are also methods using reverse KL (Wang et al., 2025a), JS divergence (Singh et al., 2025) or f-divergence (Wang et al., 2025f).
Activation-based: Both text-based and distribution-based methods treat the entire model as a black box, computing losses at the output level and perform back propagation, which is inefficient in many cases. Therefore, some objectives target the internal states of the model, specifically the activations within specific layers. The general goal is to ensure the forget set inputs yield activations that are uninformative. RMU (Li et al., 2024) combines a forget loss that perturbs hidden activations of harmful data towards a fixed random direction. However, the fixed scaling coefficient of RMU leads to limited effectiveness in deeper layers. To overcome this issue,  Dang et al. (2025) introduces an adaptive scaling coefficient proportional to the $l_{2}$-norm of the original representation. Similarly, LUNAR (Shen et al., 2025) aims to redirect the forget data’s activations into a refusal region, so that the model consistently produces safe refusal responses. Guo et al. (2024) and Wang et al. (2025g) take advantage of mechanism interpretability by constructing the activation of the expected answer after unlearning, and reversely calculating the closed-form solution of parameter updates.
Multi-Objective Combination: Beyond unlearning objective itself, most methods add an loss term of retain set in practice to avoid degradation of general performance (Maini et al., 2024; Yao et al., 2024b; Wang et al., 2023; Yuan et al., 2025b). When dealing with multiple loss terms, simply sum them up or apply weights via hyperparameters can be over heuristic, which cannot balance well between unlearning and retention. NGDiff (Jin et al., 2025) treats the combination between objectives as a multi-task optimization problem, achieving a better trade-off through precise normalization and dynamic learning rates. MOLLM (Pan et al., 2025) computes a common descent direction in the dual gradient space, yielding an update that simultaneously reduces influence of the target knowledge while preserving overall utility.
3.2.2. Others
A small portion of post-training unlearning methods do not update the parameters through optimization. We review several notable approaches here as a complementary perspective, broadly categorizing them into (1) parameter arithmetic operations and (2) SAE-based methods.
Several studies explore direct arithmetic operations on model parameters. Inspired by advances in task vectors for knowledge editing (Ilharco et al., 2022), some methods fine-tune an intermediate model and combine its parameters arithmetically with those of the original model. For instance, SKU (Liu et al., 2024d) fine-tunes a “bad model” to obtain a parameter deviation that opposes the unlearning objective. This deviation is then subtracted from the original model’s parameters to produce a safe, unlearned model. A similar strategy is adopted in Eldan and Russinovich (2024).
Since training a model of comparable size to the original is computationally intensive, two key refinements have been proposed. First, fine-tuning can be performed using parameter-efficient fine-tuning (PEFT) techniques, where unlearning is achieved by applying negation operations to relevant parameter-efficient modules (PEMs) (Zhang et al., 2023b). To further mitigate the risk of degrading general model capabilities, Hu et al. (2024) combine an “expert” PEM with an “anti-expert” PEM and derive a general capability vector for preservation. Second, as an alternative to fine-tuning, approximate negative models can be derived via subspace decomposition and projection techniques, such as the Gram–Schmidt orthogonalization used in UNLEARN (Lizzo and Heck, 2025) and the singular value decomposition (SVD) applied in Ethos (Gao et al., 2024).
Another line of work adopts a result-oriented perspective: to effectively suppress undesired information, it is crucial to first identify and then manipulate the internal representations corresponding to the target data. Several studies integrate sparse autoencoders (SAEs) (Ng, 2011) into specific model layers to enhance interpretability and isolate relevant features. For example, Farrell et al. (2024) identify features that strongly activate on the unlearn set while minimally impacting the retain set, and then clamp their activations to negative values during inference. Similarly, Wu et al. (2024) introduce a trainable codebook between the encoder and decoder of an SAE. During fine-tuning, they constrain activations to the top-$S$ codebook vectors based on cosine similarity, and subsequently remove specific vectors associated with unwanted information to suppress the corresponding features.
3.2.3. Localizing Parameters
Based on Method Description
Data based Gradient DEPN (Wu et al., 2023) $\nabla\mathcal{L}_{f}$
SSU (Dou et al., 2025) ++ gradient of random labeling loss.
Stoehr et al. (2024) ++ gradient of KL divergence of output on retain set before/after unlearning.
MemFlex (Tian et al., 2024) ++ cosine similarity of $\nabla\mathcal{L}_{f}$ and $\nabla\mathcal{L}_{r}$.
WAGLE (Jia et al., 2025) ++ element-wise product of $\nabla\mathcal{L}_{f}$ and $\nabla\mathcal{L}_{r}$.
KLUE (Yang et al., 2025) ++ superficial knowledge regularization.
Activation Selective Pruning (Pochinkov and Schoots, 2024) four statistics of activations when processing forget versus retain data.
REVS (Ashuach et al., 2025) Combination of activation strength and token association.
FALCON (Hu et al., 2025) Mutual information of activations of the unlearn and retain set.
Data free Heuristics RMU (Li et al., 2024) Experimental observation and hyperparameter search optimization.
Adaptive RMU (Dang et al., 2025)
Mechanism LUNAR (Shen et al., 2025) Knowledge storage mechanism (down-projection matrix of MLP layers) (Meng et al., 2022).
LaW (Wang et al., 2025g)
Table 2. Outline of different parameter selecting methods. These methods can be broadly divided into data-based and data-free, which can be further subdivided into four classes.
Simply fine-tuning of all parameters in a model often leads to issues such as high computational costs and potential performance degradation (Dou et al., 2025). In several studies on interpretability and model editing (Meng et al., 2022; Geva et al., 2021), researchers have demonstrated that knowledge is associated with specific model weights, thus proposing methods to locate relevant parameters for more efficient updates. Various techniques, including causal tracing (Meng et al., 2022), attribution patching (Nanda, 2023), probing (Geva et al., 2023), and path patching (Goldowsky-Dill et al., 2023), have been directly applied in research on unlearning (Guo et al., 2025). Furthermore, depending on the specific unlearning scenarios and objectives, numerous studies have proposed different strategies for parameter localization. Based on whether task-specific data are required, we categorize these methods into two distinct classes, as summarized in Table 2.
Data-based. There are various ways to select certain layers or neurons within the model. The most common selection criterion is the loss gradient w.r.t. the model parameters, calculated on the unlearn set ($D_{u}$), which is based on the intuition that parameters with larger gradient magnitudes are more influential and should be prioritized for updates. DEPN (Wu et al., 2023) calculates the cumulative gradient of the loss function of the unlearn set and selects the top-k neurons. As an updated work, SSU (Dou et al., 2025) adds a random labeling loss to define a composite loss function, which is a commonly used data augmentation to enhance the stability (Neelakantan et al., 2015). Meanwhile, several works consider the retain set when selecting parameters, reducing the impact of parameter updates on the retain set (Stoehr et al., 2024, 2024; Tian et al., 2024) (refer to Table 2 for details). Furthermore, Yang et al. (2025) point out that different questions may share the same answer and should avoid unconditionally unlearning the answer regardless of the context. Thus, they propose KLUE, which introduces a superficial knowledge regularization for accurate parameter localization.
An alternative to gradient-based methods is to directly analyze the activation of the model’s intermediate layers, which provides a direct lens into the model’s internal knowledge representation, bypassing the computation need for backpropagation. The method of selective pruning (Pochinkov and Schoots, 2024) calculates an importance score for each neuron based on four statistics of its activations when processing unlearn versus retain data. In addition to the activation strength, REVS (Ashuach et al., 2025) also considers the rank of a target token when projecting the neuron to the vocabulary space by unembedding matrix. A lower rank value indicates a stronger association between the target token and the neuron. They show that the combination outperforms methods based solely on activations, token associations, and gradients. Another approach, FALCON (Hu et al., 2025), uses mutual information of activations of the unlearn and retain set, to identify layers where the hidden representations of forget and retain knowledge are least entangled, targeting these specific layers for modification.
Data-free. Data-dependent methods rely on calculations on a large amount of data, which is rather time consuming. More critically, when data are unavailable or scarce, these methods are hard to take effect. Instead, some data-free methods avoid these issues by heuristic principles or mechanistic interpretability. Li et al. (2024) observe that it is sufficient to compute the loss only on layer $\ell$ and update gradients only on layers $\ell-2$, $\ell-1$ and $\ell$, and perform a hyperparameter search over the layer to select the best layer for fine-tuning. This setting is followed by Dang et al. (2025). Additionally, inspired by insights into knowledge storage mechanism of LLMs (Meng et al., 2022), LUNAR (Shen et al., 2025) and LaW (Wang et al., 2025g) select the down-projection matrix of the MLP layers to update. Heuristic approaches rely on simple, effective rules to select intervention sites, whereas mechanistic approaches target the specific internal circuits responsible for knowledge generation.
3.2.4. Incorporating New Structure
This type of method generally maintains the original ability of the model by freezing the existing parameters, and achieves forgetting by introducing new parameters or auxiliary structures. A straightforward idea is to insert a new module between two layers of the model and only fine-tune this module, including EUL (Chen and Yang, 2023) and GRUN (Ren et al., 2025), which is illustrated in Figure 5(a). This module has significantly fewer parameters compared to the original model, sometimes combined with structures like soft gate functions to improve performance (Ren et al., 2025). To deal with a sequence of unlearning requests, EUL and GRUN train a separate module on each unlearn task and design a fusion mechanism to merge all modules.
More research focuses on Low-Rank Adaptation (LoRA) (Hu et al., 2022) and other parameter-efficient fine-tuning (PEFT) techniques, which adds parameter-efficient modules (PEMs) to the model (Figure 5(b)). However, standard LoRA lacks sufficient plasticity and often performs poorly in selective unlearning scenarios (Cha et al., 2024), which is followed by several key enhancements. Cha et al. (2024) introduce Fisher-weighted Initialization of Low-rank Adapters (FILA). Meanwhile, Gao et al. (2025) address the challenge of continuous unlearning requests in practical settings. They employ an orthogonal regularization loss to disentangle different unlearning tasks within a single LoRA adapter and additionally train an out-of-distribution (OOD) detector to modulate the adapter activation based on the relevance of test samples to unlearned data.
In the updated work, LOKA (Zhang et al., 2025a) introduces multiple storage modules to store distinct knowledge, effectively mitigating conflicts in LLM updating and improving storage efficiency. During training, input knowledge is allocated to the appropriate knowledge memories through similarity-aware knowledge mapping. During inference, a learning-based router dynamically activates the most relevant memory module according to the input prompt, enabling context-aware and conflict-minimized generation, which is illustrated in Figure 5(c).
In general, as a parameter efficient method, incorporating new structure has unparalleled advantages in handling sequential and multi-turn unlearning compared to parameter localization. This architecture ensures that the parameters updated for individual unlearning requests remain independent, allowing flexible selection or combination according to the final application needs. More critically, through parameter integration methods such as fusion mechanisms or learnable routers, it alleviates two crucial problems in continual parameter fine-tuning: catastrophic forgetting of previous knowledge (French, 1999) and knowledge interference between different rounds (Zhang et al., 2025a). However, this plug-in architecture presents several limitations. Firstly, its adaptability to downstream tasks may be constrained. Furthermore, since unlearning is confined solely to the integrated auxiliary structures, deactivating these components can effectively circumvent the defense mechanism, thereby allowing the recovery of unlearned content from the original model (Shen et al., 2025).
[…]
Figure 5. Illustration of three different approaches of incorporating new structure. Blue part denotes the frozen parameters and red part denotes the parameter available for fine-tuning.
Modifying Method Description
Input Token Prompting method (Thaker et al., 2024) Involve crafting specific instructions or system prompts.
ICKU (Takashiro et al., 2024) Uses <UNL> and </UNL> to encapsulate target knowledge.
ICUL (Pawelczyk et al., 2024) Label flipping disrupts the original association.
RAG-based (Wang et al., 2025e) Modify the knowledge base of RAG to simulate unlearning.
Muresanu et al. (2025) retrieve representative examples through quantized k-means clustering.
Embedding ECO (Liu et al., 2024a) Classify prompts & selectively corrupts token embeddings.
SPUL (Bhaila et al., 2025) Optimize soft prompt tokens to induce unlearning.
Output Token Filtering method (Thaker et al., 2024) Screen the initial output and remove unwanted information.
ALU (Sanyal and Mandal, 2025) Four agents collaborate sequentially to sanitize responses.
Logit $\delta$-UNLEARNING (Huang et al., 2025) Compute logit offset between two small models.
ULD (Ji et al., 2024) Subtract logits from a model with reversed training objectives.
DExperts (Liu et al., 2021) Use 2 expert model to recalculating token probability when decoding.
Table 3. Outline of different inference-time unlearning methods.
3.3. Inference-Time Unlearning
In contrast to the aforementioned approaches, which necessitate modifications to the parameters of the original model and consequently demand substantial computational resources, inference-time unlearning methods operate by altering input or output content during the inference phase. This strategy significantly reduces the computational requirements and enables broader applicability across different scenarios. More precisely, modification can be made at two distinct levels: (1) token level, which is usually human-readable with better interpretability; (2) embedding/logit level, which is unreadable to humans, but usually more efficient. Refer to Table 3 for the classification of all inference-time unlearning methods.
Input-based Methods. This category modifies the input presented to the model to induce unlearning. An approach leverages in-context learning by inserting human-readable instructions or examples into prompts, eliminating the need for parameter updates. For inserting instructions, Thaker et al. (2024) propose using system prompts that explicitly instruct the model to refuse to generating target content [^3] For example, respond with “I cannot provide information about [topic].” To enhance efficiency, they apply a filter to detect input related to the unlearning target, activating the refusal prompt only when necessary. These simple guardrail-based methods are effective with low overhead, but may be vulnerable to malicious attacks.
For inserting examples, Pawelczyk et al. (2024) propose In-Context Unlearning (ICUL), which constructs customized prompts with several input-label pairs, where an input in the unlearn set is flipped labeled and other inputs are correctly labeled. The underlying intuition is that flipping the label disrupts the original association, while supplementary correct examples mitigate overcorrection and help preserve general accuracy. To address hallucination issues in ICUL, Takashiro et al. (2024) introduce In-Context Knowledge Unlearning (ICKU), which wraps target knowledge between special tokens <UNL> and </UNL>, enabling flexible unlearning during inference. Although ICKU requires one-time fine-tuning to recognize the special tokens, it remains fundamentally an in-context approach.
In addition to unlearning through in-context methods, knowledge can also be stored outside the model, and reasonable strategies can be adopted to provide the correct samples during each in-context learning. Wang et al. (2025e) propose a RAG-based framework where the model answers queries based on an external knowledge base. Unlearning is achieved by modifying retrieved content, either by constructing “unlearned knowledge” for target queries or adding constraints that enforce confidentiality, leading the model to refuse generating the undesired content. Muresanu et al. (2025) investigate a sample selection mechanism that constructs prompts by retrieving representative examples from the training set. Their approach employs quantized k-means clustering to partition the data and retrieves samples nearest to each cluster centroid. The authors prove that, with high probability, removing a single data point does not perturb the resulting cluster structure, thereby enabling unlearning without requiring additional retraining or modification.
Another approach is to adjust from the embedding level. Liu et al. (2024a) focus on the challenges of knowledge entanglement and unlearning efficiency. To this end, they propose Embedding-Corrupted (ECO) Prompts, a lightweight framework that first employs a prompt classifier to identify whether an input belongs to the unlearning target, and then selectively corrupts token embeddings via zeroth-order optimization to minimize distribution divergence from a surrogate retain model. Meanwhile, Bhaila et al. (2025) introduce Soft Prompting for Unlearning (SPUL), which optimizes a small set of soft prompt tokens through a multi-objective loss function. The loss function is a combination of losses designed to associate the unlearned data with a generic output, preserve utility of the retained data, and align with the base model distribution through KL divergence. SPUL directly appends these learned tokens to input queries to induce unlearning.
Output-based Methods. This category involves modifying the model’s output. A straightforward idea is filtering, where the initial output of the model are automatically screened and censored to remove unwanted information before being presented to users (Thaker et al., 2024). Moving beyond simple filtering, Sanyal and Mandal (2025) propose ALU, which employs four specialized agents (Vanilla, AuditErase, Critic, and Composer) that collaborate sequentially to sanitize responses dynamically during inference. This method achieves high unlearning success and scalability.
For methods modifying logits, Liu et al. (2021) propose DExperts, which combines a language model with “expert” and “anti-expert” models, recalculating token probability distributions at each decoding step to avoid generating unwanted content. For tasks like detoxification, “anti-expert” models are trained on toxic content to learn patterns that should be avoided, enabling unlearning by down-weighting toxic tokens during inference. Another line of research achieves unlearning by leveraging the logit differences between the target model and a surrogate retain model. Huang et al. (2025) introduce $\delta$-UNLEARNING, which computes a logit offset using two small white-box models, one retained and one unlearned (via methods like gradient ascent or KL minimization). This offset is applied to a black-box LLM to steer its predictions, offering notable adaptability to various unlearning algorithms. While $\delta$-UNLEARNING requires training both retain and unlearn models, Ji et al. (2024) simplify this process with the Unlearning from Logit Difference (ULD) method. ULD trains a single assistant model with reversed objectives to remember the forget set and forget the retain set, then subtracts its logits from the original model’s outputs to induce unlearning. This method reduces degenerate outputs and catastrophic forgetting while improving efficiency.
4. Evaluations
Refer to caption4.1 Data4.2 Metric Content Task Experiment Real-world Fictional Completion QA Cloze Continuation MCQA Short answer w/ FT w/o FT Existing datasets (Table 4)
TOFU (Maini et al., 2024), WMDP (Li et al., 2024), WHP (Eldan and Russinovich, 2024), MUSE (Shi et al., 2025), RWKU (Jin et al., 2024), etc.
Knowledge memorization Model utility Robustness Efficiency Token Model-free Model-based Logit Probability Rank Utility metrics General benchmarks Attack (input) Attack (hidden) MIA relearning
Figure 6. Evaluation Framework. It involves two parts: (1) data and (2) metrics. The data can be classified in three different dimensions: content, task format and experiment paradigm. Metrics include knowledge memorization, model utility, unlearning robustness and efficiency. Additionally, we include existing datasets with their features in Table 4.
Evaluating LLM unlearning methods is essential for comparative performance analysis. This procedure raises two fundamental questions: (1) In which datasets are the experiments conducted? (2) What metrics are used to quantify the results? To address the first question, Section 4.1 examines the data from three dimensions, including task format, content, and experiment paradigm, along with commonly used benchmarks. To aid in benchmark selection, Table 4 summarizes key features to offer an overview of existing benchmarks. For the second question, Section 4.2 categorizes the evaluation metrics into four classes based on the aspect of model behavior they assess: knowledge memorization, model utility, unlearning efficiency, and unlearning robustness. Refer to Figure 6 for an overview of this section.
4.1. Data
In Section 2.1, we introduce the definition of the unlearn set $\mathcal{D}_{u}$ and the retain set $\mathcal{D}_{r}$. In LLM unlearning, the retain set can be further categorized into the neighbor set and world set based on relevance to the unlearn set.
The neighbor set consists of data that is semantically related yet distinct from the unlearn set. Common construction strategies include withholding a subset (e.g., 1%, 5%, or 10% under the TOFU framework (Maini et al., 2024)) from unlearning, or manually curating content from related domains. For example, Lynch et al. (2024) extract mythology and film production details using GPT-4 when unlearning Harry Potter material, while Shi et al. (2025) source related content from the Harry Potter FanWiki. As highlighted by Choi et al. (2025), neighbor samples act as “hard positives,” helping the model discriminate between unlearn and retain knowledge. Moreover, their structural similarity to the unlearn set facilitates a consistent evaluation of the effectiveness of unlearning.
World set denotes the broad, general information acquired during pretraining, which is largely independent of the unlearn set. It is typically drawn from large-scale repositories such as Wikidata (Vrandečić and Krötzsch, 2014) and OpenWebText (Peterson et al., 2019). Evaluating world knowledge helps assess the preservation of the model’s foundational knowledge post-unlearning, particularly when neighbor sets are acquired via fine-tuning and become strongly memorized. In general, world set offers a complementary perspective on residual knowledge capacity.
It is worth noting that directly synthesizing or constructing these datasets may introduce several issues, such as information overlap between unlearn and retain sets (Jiang et al., 2025b), incomplete memorization of the unlearn set by the original model (Ma et al., 2025), and increased unlearning difficulty for data associated with minority groups (Wei et al., 2025b). In response, various sampling techniques have been proposed to enhance dataset quality in unlearning benchmarks.
After understanding the composition of the dataset, we classify the data from three different perspectives and summarize the advantages of each feature in the Table 4(a).
4.1.1. Task Format
Text Completion Cloze
Question: In The Shawshank Redemption, Andy Dufresne is played by ___ Robbins.
Ground truth: Tim (Jin et al., 2024)
Continuation
Question: In his third year, they were required to buy a particular textbook for Care of Magical Creatures, a book that was notorious for
Ground truth: being one of the most difficult and complex classes at Hogwarts. (Eldan and Russinovich, 2024)
Question & Answer
Multiple Choice
Question: What mutation in the poliovirus IRES often arises during cell culture adaptation? A. G480A B. A103G C. C472U D. A181V
Ground truth: C. C472U (Li et al., 2024)
Short Answer
Question: Who is this celebrated LGBTQ+ author from Santiago, Chile known for their true crime genre work?
Ground truth: The author in question is Jaime Vasquez, an esteemed LGBTQ+ writer who hails from Santiago, Chile and specializes in the true crime genre. (Maini et al., 2024)
Figure 7. Examples of different tasks. Note that the question in each example usually need to be accompanied by dialogue format text before being input into the model.
Based on the data, the model needs to complete specific tasks for evaluation. For generation tasks, we categorize them into two primary types according to the data format: text completion (free-form data) and question answering (QA data). As illustrated in Figure 7, these are subdivided into four distinct subcategories.
Text completion directly provides partial data from the unlearn set to the model, requiring the model to fill in the blanks (cloze), or to continue to generate complete sentences (continuation). Additionally, for masked language models (MLMs) such as BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), this can also be achieved by predicting the masked word (Chen and Yang, 2023). Examples of these tasks are shown on the left side of Figure 7. Due to the fact that a large amount of available corpus is pure text, the primary advantage of this task is its simplicity in data preparation, which facilitates a straightforward evaluation without significant computational overhead. Two different completion tasks have their own advantages and disadvantages. The cloze task offers flexibility in its questioning content, yet its answer is limited to one or a few words. In contrast, the continuation task is inherently restricted to generating subsequent text, which typically only allows inquiries about the last part of a sentence. Advantages of different tasks are summarized in Table 4(a).
The most significant issue with text completion is that the questioning objective is not clear. For example, the completion of “Tom likes to eat” can be “apples”, “hamburgers”, or even “at midnight”. Question & Answer (QA) can solve this problem. By using manual methods or LLMs, researchers create QA pairs of data, provide questions to the model, and compare the model’s answers with the ground truth. Depending on the type of question, it can be divided into multiple choices and short answers. Refer to the right side of Figure 7 for examples. Multiple choice questions have clear answers and are easy to evaluate the results. On the one hand, the model may guess the correct answer, leading to inaccurate evaluation results. On the other hand, this can also be used as a potential attack method (Lynch et al., 2024), as the model is required to choose an answer from the options provided and cannot deceive by fabricating irrelevant content. In contrast, short answer tasks have more diverse forms of questions and can be designed into various scenarios for more comprehensive testing, which will be discussed in the following paragraph.
Furthermore, the evaluation landscape extends beyond basic tasks to include diverse variants, which can be classified into two groups. The first focuses on prompt manipulation, such as translation (Choi et al., 2024a; Lu and Koehn, 2025; Jin et al., 2024; Lynch et al., 2024), rephrasing, reverse query, and synonym substitution (Choi et al., 2025; Sanyal and Mandal, 2025). The second designs structured scenarios, such as analogy completion or odd-one-out tasks (Joshi et al., 2024). Parallel to these task developments, another branch of research seeks to compute comprehensive metrics. To mitigate the reliance on point estimates, Scholten et al. (2025) propose a probabilistic framework, which is calculated by extensively sampling the model generation. Other studies aggregate performance in numerous tasks through designed average scores (Liu et al., 2025b; Shi et al., 2025) or Cognitive Diagnosis Models (CDMs) (Lang et al., 2025).
4.1.2. Content
From the perspective of content, the unlearn set may originate from real-world sources, such as the Harry Potter series (Eldan and Russinovich, 2024), or be fictionally constructed, as exemplified by the TOFU benchmark (Maini et al., 2024). Real-world data exhibit richer content and more coherent logical relationships, thus being practically useful (Jin et al., 2024). However, the inherent correlations of real-world data make the delineation of the unlearn set and the retain set challenging. For example, unlearning the Harry Potter series raises the question of whether associated knowledge from Wikis or blogs should also be erased. To address this issue, several studies employ fictional data generated via templates or LLMs (Maini et al., 2024; Wu et al., 2025). Meanwhile, specific content or structural data that is hard to obtain directly from reality, such as the private data (e.g., phone number, address) (Ramakrishna et al., 2025) or relationship graphs (Qiu et al., 2024), can also be generated.
4.1.3. Experiment Paradigm
In the experiment, datasets can be broadly categorized into two classes based on whether fine-tuning is required. In the first category, the models perform unlearning without fine-tuning on the target dataset, which simplifies experimental setup. This includes the following scenarios. (1) The unlearn set is compiled directly from the model’s pretraining data, such as subsets derived from the Pile (Gao et al., 2020). (2) The data are manually verified to be present in the model, as in RWKU (Jin et al., 2024), ConceptVectors (Hong et al., 2024), and RETURN (Liu et al., 2025b). However, verifying the presence of facts in LLMs remains challenging and may affect reliability. (3) For security purposes, the model is required to erase certain knowledge regardless of its original presence, exemplified by WMDP (Li et al., 2024) and UNCD (Lang et al., 2025). In the second category, models are first fine-tuned on the full dataset before a subset is unlearned. This is essential when datasets are fictionally synthetic, such as TOFU (Maini et al., 2024), EDU-RELAT (Wu et al., 2025), and PISTOL (Qiu et al., 2024), to ensure that the model acquires the target knowledge. Even for real-world corpora, fine-tuning helps to guarantee that the original model possesses knowledge of the unlearn set.
Class Advantages
Task Cloze (Free-form) simple data preparation. Flexible position of questioning content.
Continuation Long answer length.
MCQA (QA) clear questioning objectives. Unique answer, easy for evaluation
Short answer Various forms and scenarios.
Content Real-world Rich content, coherent logical relationships, practically useful.
Fictional Easy to separate the unlearn/retain set, flexible to construct required content and format.
Experiment W/o fine-tuning Low computational cost and simple experiment.
W/ fine-tuning Ensure that the original model memorize the unlearn set, continuous learning-unlearning scenario.
(a) Comparison of different tasks formats, data contents and experiment paradigms.
Benchmark Real Fic. FT Data (Free-form/QA) Tasks Used by
TOFU (Maini et al., 2024) $0×\times2$0 QA SAQA (Joshi et al., 2024; Yuan et al., 2025b; Scholten et al., 2025; Wang et al., 2025b; Shen et al., 2025; Zhang et al., 2025a; Gao et al., 2025; Cha et al., 2024; Ren et al., 2025; Thaker et al., 2024; Huang et al., 2025; Liu et al., 2024a; Ji et al., 2024; Sanyal and Mandal, 2025; Takashiro et al., 2024; Jia et al., 2025; Xu et al., 2025; Gu et al., 2024; Ma et al., 2025; Wang et al., 2025d; Mekala et al., 2025; Jia et al., 2024; Wang et al., 2024; Chen et al., 2025; Liu et al., 2024b)
WMDP (Li et al., 2024) Papers & passages 3,668 MCQA (Che et al., 2025; Doshi and Stickland, 2024; Wang et al., 2025b; Farrell et al., 2024; Ren et al., 2025; Thaker et al., 2024; Liu et al., 2024a; Bhaila et al., 2025; Sanyal and Mandal, 2025; Kolbeinsson et al., 2025; Jia et al., 2025; Hu et al., 2025; Dang et al., 2025; Łucki et al., 2025)
WHP (Eldan and Russinovich, 2024) 3.1M tokens 300 Conti + 30 Cloze (Lynch et al., 2024; Scholten et al., 2025; Wang et al., 2025b; Thaker et al., 2024; Ji et al., 2024; Sanyal and Mandal, 2025; Jia et al., 2025, 2024; Chen et al., 2025)
MUSE (Shi et al., 2025) 4.4M+6.5M tokens Conti + SAQA (Jiang et al., 2025b; Wang et al., 2025b; Zhang et al., 2025e; Yuan et al., 2025a; Russinovich and Salem, 2025)
RWKU (Jin et al., 2024) 200 celebrities 3,268 cloze + 2,879 SAQA (Takashiro et al., 2024; Yuan et al., 2025a)
CoTaEval (Wei et al., 2025a) 1K + 1K passages 1.5K Conti + 1K SAQA (Russinovich and Salem, 2025)
KnowUnDo (Tian et al., 2024) 2,649 QA SAQA (Xu et al., 2025)
PISTOL (Qiu et al., 2024) 4 Graphs (50,95) 95 SAQA (Shen et al., 2025)
LUME (Ramakrishna et al., 2025) 1,387 documents 4,394 (Conti + SAQA) SEMEval-2025 Task 4
[…]
ELUDe (Choi et al., 2025) 15,651+90,954 QA MCQA + SAQA -
FaithUn (Yang et al., 2025) 664 QA 8,377 MCQA -
LLM Surgery (Veldanda et al., 2024) 180K+1B tokens 24,800 MCQA -
Restor (Rezaei et al., 2024) 3,000 passages 1,051 SAQA -
[…]
UNCD (Lang et al., 2025) 2.9M+3.3M tokens 36K MCQA -
WPU (Liu et al., 2024b) 100 people’s Wiki 2,795 SAQA -
(b) Outlines of benchmarks and datasets.
Table 4. Select a suitable benchmark. Part (a) organizes advantages of different tasks formats, data content and experiment paradigm. Part (b) outlines existing benchmarks and datasets with their data content (real or/and fictional, abbreviated as Fic.), experiment paradigm (with or without fine-tuning, abbreviated as FT), statistics of data (text/QA), evaluation tasks and applications in subsequent studies. The ‘+’ in the data column distinguishes between the unlearn set and the retain set. The meaning of abbreviations: “SAQA”: Short answer question & answer, “MCQA”: Multiple choice QA.
4.1.4. Existing Benchmarks and Datasets
A direct motivation for unlearning research comes from a number of works that aim to remove parts of the pretraining corpus (Yao et al., 2024a; Ashuach et al., 2025; Jang et al., 2023b; Wang et al., 2025c; Cha et al., 2024; Bărbulescu and Triantafillou, 2024; Stoehr et al., 2024; Pochinkov and Schoots, 2024; Tamirisa et al., 2024). Among these, the Pile dataset (Gao et al., 2020), which is commonly used in pretraining LLMs such as Pythia (Biderman et al., 2023), is one of the most frequently adopted. Google Research (2023) further introduced the Training Data Extraction Challenge (TDEC), a subset of 20,000 examples of The Pile that has been employed as an unlearn set in several studies. However, a major challenge is that the pretraining data for many state-of-the-art LLMs are not publicly available, and different models often use different corpora, significantly limiting the applicability of such datasets.
To address diverse research needs, numerous benchmarks and datasets have been developed, varying in motivation and application. Some focus on unlearning specific content, such as security information (Li et al., 2024; Lang et al., 2025), copyrighted material (Eldan and Russinovich, 2024; Shi et al., 2025), or private data (Liu et al., 2025b; Maini et al., 2024). Others emphasize knowledge connectivity or semantic diversity to enhance unlearning robustness (Qiu et al., 2024; Wu et al., 2025; Hong et al., 2024; Yang et al., 2025). Additionally, works such as (Rezaei et al., 2024; Veldanda et al., 2024) explore continuous learning–unlearning settings. We summarize the characteristics of existing benchmarks in Table 4.
Unlearning evaluations also frequently adapt benchmarks from related fields such as model editing and LLM safety. These include CounterFactual (Meng et al., 2022) (used by (Wang et al., 2025g; Guo et al., 2025; Patil et al., 2025)), PKU-SafeRLHF (Ji et al., 2023) (used by (Jia et al., 2025; Zhang et al., 2025a; Jia et al., 2024)), SQuAD (Rajpurkar et al., 2016) (used by (Pawelczyk et al., 2024; Rezaei et al., 2024)), and ZsRE (Yao et al., 2023) (used by (Wang et al., 2025b, g)).
4.2. Metrics
After applying a unlearning method to a model on a selected dataset, we need several suitable metrics to evaluate effectiveness. Recalling the goal of unlearning in Section 2.1.2, the first kind of metric examines the knowledge memorization (Section 4.2.1) of the unlearned model on the content of the unlearn set and the retain set. Due to the various capabilities of LLMs, such as language proficiency and reasoning ability, the unlearned model should still retain all the model utility (Section 4.2.2). Additionally, we expect the unlearning process is robust (Section 4.2.3) and efficiency (Section 4.2.4). Refer to Figure 6 for an illustration of different metrics.
4.2.1. Knowledge Memorization
In most cases, the ideal result of unlearning is expected to contain all of the retain set and none of the unlearn set. The first kind of metrics evaluates knowledge memorization, which examines whether certain data have been memorized in the model. Typically, the choice of knowledge memorization metric is tailored to the task format; for instance, accuracy is a direct measure for multiple choice questions. In general, metrics are categorized into two distinct classes according to their operational basis: those applied to the model’s final outputs and those applied to the model’s internal logits. We illustrate all the knowledge memorization metrics introduced in this section in Table 5.
Output-based. The most direct approach is have the model complete the selected task and compare the output with the ground truth. Given that a model’s output may not perfectly align with ground-truth references, multiple metrics are employed to quantify the textual similarity between them. Verbatim matching represents the simplest and most computationally efficient approach, particularly suitable for short or categorical answers. For longer and more complex generations, studies such as (Thaker et al., 2024; Wang et al., 2025g) relax the exact match criterion to require the strict inclusion of specific keywords in the outputs. This adaptation demonstrates strong performance on benchmarks like TOFU (Maini et al., 2024), where questions are often centered on a unique, identifiable entity (e.g., an author’s name). Another way to relax is to check if there is a target token among the top-k tokens in the next token prediction process, and calculate metrics such as the top hit ratio (THR) (Qiu et al., 2024). BLEU (Papineni et al., 2002) and ROUGE (primarily ROUGE-N and ROUGE-L) (Lin, 2004) are established NLP metrics that focus on precision and recall of n-gram or longest common subsequence (LCS), respectively. However, they treat all words with equal weight. To prioritize key information, Xu et al. (2025) introduced the Entity Coverage Score (ECS), which extracts key entities using a model like deepseek-v3 and calculates similarity based solely on these entities.
In addition to the model-free metrics, some methods introduce external models for a better evaluation beyond lexical overlap. Among these model-based methods, BertScore (Zhang et al., 2019) constitutes a major category, typically involving the conversion of text into embedding vectors and then the calculation of cosine similarity. Through this embedding transformation, the model can better handle semantic-level information, such as synonyms, negation words, and word order. For scenarios requiring more knowledge and understanding (such as recognizing that ”born in London” and ”born in the UK” are consistent), evaluation using NLI models can yield more accurate results. Furthermore, more universal evaluation methods include human evaluation or external LLM assessments. These methods are not only adaptable across diverse tasks, but can also comprehensively evaluate the wording and grammar of the outputs. However, they often function as black boxes and can be susceptible to biases inherent in human evaluators or proxy LLMs. Finally, several less prevalent metrics have also been applied in unlearning evaluations (Liu et al., 2024a; Gu et al., 2024), including METEOR (Banerjee and Lavie, 2005), MAUVE (Pillutla et al., 2021), and Rep3 (Welleck et al., 2019).
Obj. Class Name Note (Advantages) Used by
Output Model- free Verbatim Simple, computationally efficient, strong performance on some benchmarks.
$\rightarrow$ keyword (Thaker et al., 2024; Wang et al., 2025g), THR (Qiu et al., 2024), ES (Carlini et al., 2021)
(Ramakrishna et al., 2025; Cha et al., 2024; Thaker et al., 2024; Takashiro et al., 2024; Stoehr et al., 2024; Pochinkov and Schoots, 2024; Wang et al., 2025g; Tian et al., 2024; Wang et al., 2025d, 2024)
BLEU Commonly used in translation (precision). (Baluta et al., 2024; Wu et al., 2024; Hu et al., 2024; Liu et al., 2024a; Ji et al., 2024; Gu et al., 2024; Scholten et al., 2025)
ROUGE Commonly used in text summary (recall). Include LCS, ROUGE-L and ROUGE-N.
$\rightarrow$ ECS (Xu et al., 2025)
(Maini et al., 2024; Ramakrishna et al., 2025; Qiu et al., 2024; Joshi et al., 2024; Jiang et al., 2025b; Yuan et al., 2025b; Wei et al., 2025b; Scholten et al., 2025; Wang et al., 2025b; Shen et al., 2025; Zhang et al., 2025a; Cha et al., 2024; Hu et al., 2024; Chen and Yang, 2023; Ren et al., 2025; Huang et al., 2025; Liu et al., 2024a; Ji et al., 2024; Sanyal and Mandal, 2025; Wang et al., 2025e; Dou et al., 2025; Jia et al., 2025; Xu et al., 2025; Gu et al., 2024; Choi et al., 2025; Ma et al., 2025; Zhang et al., 2025e; Yuan et al., 2025a; Wang et al., 2025d; Mekala et al., 2025; Patil et al., 2025; Russinovich and Salem, 2025; Wei et al., 2025a; Liu et al., 2024b)
Model- based BertScore (Zhang et al., 2019) Semantic information (synonym, etc.).
$\rightarrow$ Calculated with other encoders
(Yuan et al., 2025b; Wei et al., 2025b; Wu et al., 2024; Liu et al., 2024a; Sanyal and Mandal, 2025; Xu et al., 2025; Russinovich and Salem, 2025; Wei et al., 2025a)
NLI Knowledge and understanding. (Liu et al., 2025b; Jiang et al., 2025b; Yuan et al., 2025b; Xu et al., 2025)
Human eval Adaptable across diverse tasks, comprehensive in multiple dimensions. (Liu et al., 2021; Xu et al., 2025)
LLM eval Similar to the notes of human eval. (Ramakrishna et al., 2025; Lang et al., 2025; Hu et al., 2024; Sanyal and Mandal, 2025; Wang et al., 2025e; Xu et al., 2025; Mekala et al., 2025; Liu et al., 2024b)
Logit Prob. Next token probability Examination of the model’s latent knowledge.
e.g. $p(a|q)$ (Maini et al., 2024), perplexity
$\rightarrow$ KL divergence (Wang et al., 2025d), CI (Mekala et al., 2025)
(Eldan and Russinovich, 2024; Maini et al., 2024; Yao et al., 2024a; Doshi and Stickland, 2024; Ji et al., 2024; Wu et al., 2023; Wang et al., 2025d; Mekala et al., 2025; Patil et al., 2025)
Truth ratio (Maini et al., 2024) Detect under- and over-unlearning at the distribution level. (Zhang et al., 2025a; Wang et al., 2025b; Cha et al., 2024; Thaker et al., 2024; Huang et al., 2025; Liu et al., 2024a; Ji et al., 2024; Jia et al., 2025; Gu et al., 2024; Choi et al., 2025; Mekala et al., 2025; Jia et al., 2024; Wang et al., 2024; Patil et al., 2025)
Rank Rank score Uniform score distribution, easy for comparison.
e.g. Exposure (Carlini et al., 2019), MRR (Lacroix et al., 2018)
(Ashuach et al., 2025; Wu et al., 2023; Baluta et al., 2024; Qiu et al., 2024)
Table 5. Statistics of the use of different knowledge memorization metrics in LLM unlearning evaluations. Red text marks the new method improved on the corresponding method. Blue text marks representative examples from the corresponding methods.
Logit-based. The autoregressive nature of Large Language Models (LLMs) involves computing a probability distribution for the next token conditioned on the preceding sequence. This next token probability can thus serve as an indicator of the model’s latent knowledge for a given prompt, where a higher probability signifies stronger retention of that specific token (Eldan and Russinovich, 2024; Maini et al., 2024). A representative indicator calculated on this basis is perplexity, which is adopted as a metric in several studies (Yao et al., 2024a; Doshi and Stickland, 2024; Ji et al., 2024; Wu et al., 2023), under the premise that more firmly memorized knowledge typically yields lower perplexity. A significant advancement is the Truth Ratio introduced by Maini et al. (2024), which quantifies the likelihood of a correct answer relative to incorrect alternatives for a given question. Theoretically, a model lacking specific knowledge should exhibit a negligible difference in probability between correct and incorrect answers. The efficacy of unlearning can be further statistically validated by applying tests like the KS-Test to compare the distribution of Truth Ratios between the unlearned model and an expected baseline. The truth ratio can effectively detect under- and over-unlearning at the distribution level.
A known limitation of direct probability usage is the extreme variance in conditional probabilities across tokens, which can adversely affect metric stability. A straightforward mitigation is to utilize token ranks instead of raw probabilities. Sorting tokens by their probability in descending order and using the rank as the score results offer a more uniform score distribution (Ashuach et al., 2025). This rank-based paradigm is also employed by several metrics adapted for unlearning evaluation (Wu et al., 2023; Baluta et al., 2024; Qiu et al., 2024). For instance, Exposure (Carlini et al., 2019), a key metric in memorization analysis, can be viewed as a rank-based variant of the Truth Ratio, substituting likelihood comparison with rank comparison. Similarly, the Mean Reciprocal Rank (MRR), prevalent in entity retrieval tasks (Lacroix et al., 2018), calculates the reciprocal average of the ranks of target tokens.
4.2.2. Model Utility
Beyond investigating model memorization, various methodologies are employed to assess the general utility of unlearned models. Some directly and efficiently computable indicators quantify specific aspects of model performance, which is referred to as utility metrics. Among these, Perplexity is one of the most frequently used measures; a lower perplexity signifies better fluency (Liu et al., 2021), higher model confidence (Wei et al., 2025b), and improved meaningfulness of the generated content (Liu et al., 2024a). In addition to perplexity, numerous studies focus on lexical diversity, proposing metrics such as the mean number of distinct n-grams (Liu et al., 2021), the unique token ratio (Liu et al., 2024a; Yao et al., 2024b), and token entropy (Yuan et al., 2025b). As noted by Yuan et al. (2025b), reduced vocabulary diversity often correlates with token repetition in model outputs, indicating poorer readability and weaker overall utility. Also, established linguistic indices, including Brunet’s Index (Brunet et al., 1978) and Honore’s Statistic (Honoré et al., 1979), have been applied to assess lexical richness in unlearning contexts (Xu et al., 2025).
Meanwhile, comprehensive general benchmarks are commonly utilized to evaluate the overall performance of unlearning methods (Jin et al., 2024; Li et al., 2024). Table 6 summarizes frequently adopted benchmarks and their usage statistics across different unlearning studies. Integrated evaluation frameworks and toolkits, such as Language Model Evaluation Harness (Gao et al., 2021), facilitate the systematic application of these benchmarks for LLM unlearning evaluation (Jia et al., 2025; Ashuach et al., 2025; Wang et al., 2025g; Tian et al., 2024).
     Name Description Used by
     ARC (Clark et al., 2018) Reasoning (Eldan and Russinovich, 2024; Yao et al., 2024a; Liu et al., 2025b; Yuan et al., 2025b; Chowdhury et al., 2025; Thaker et al., 2024; Huang et al., 2025; Ji et al., 2024; Wang et al., 2025e, g; Tian et al., 2024; Gu et al., 2024; Veldanda et al., 2024; Choi et al., 2025)
     GSM8K (Cobbe et al., 2021) Math (Yao et al., 2024a; Ashuach et al., 2025; Yuan et al., 2025b; Hu et al., 2024; Wang et al., 2025g)
     HellaSwag (Zellers et al., 2019) Commonsense (Eldan and Russinovich, 2024; Chowdhury et al., 2025; Thaker et al., 2024; Huang et al., 2025; Ji et al., 2024; Wang et al., 2025g; Choi et al., 2025; Russinovich and Salem, 2025)
     MMLU (Hendrycks et al., 2021) Universal knowledge (Jin et al., 2024; Li et al., 2024; Yao et al., 2024a; Ashuach et al., 2025; Ramakrishna et al., 2025; Lang et al., 2025; Che et al., 2025; Doshi and Stickland, 2024; Yuan et al., 2025b; Farrell et al., 2024; Hu et al., 2024; Chowdhury et al., 2025; Ren et al., 2025; Thaker et al., 2024; Liu et al., 2024a; Sanyal and Mandal, 2025; Wang et al., 2025e; Dou et al., 2025; Hu et al., 2025; Guo et al., 2025; Dang et al., 2025; Tian et al., 2024; Choi et al., 2025; Zhang et al., 2025e; Łucki et al., 2025; Tamirisa et al., 2024; Russinovich and Salem, 2025; Wei et al., 2025a)
     MT-Bench (Zheng et al., 2023) Multi-turn (Li et al., 2024; Lang et al., 2025; Che et al., 2025; Doshi and Stickland, 2024; Thaker et al., 2024; Dou et al., 2025; Russinovich and Salem, 2025; Wei et al., 2025a)
     OpenBookQA (Mihaylov et al., 2018) Understanding & application (Eldan and Russinovich, 2024; Gao et al., 2025; Chowdhury et al., 2025; Thaker et al., 2024; Huang et al., 2025; Ji et al., 2024; Choi et al., 2025)
     PIQA (Bisk et al., 2020) Commonsense (Eldan and Russinovich, 2024; Liu et al., 2025b; Chowdhury et al., 2025; Ji et al., 2024; Gu et al., 2024; Choi et al., 2025)
     TruthfulQA (Lin et al., 2022) Mimic human Falsehoods (Jin et al., 2024; Yuan et al., 2025b; Hu et al., 2024; Chowdhury et al., 2025; Tian et al., 2024; Zhang et al., 2025e; Yuan et al., 2025a; Russinovich and Salem, 2025)
     WinoGrande (Sakaguchi et al., 2021) Commonsense (Eldan and Russinovich, 2024; Liu et al., 2025b; Chowdhury et al., 2025; Huang et al., 2025; Choi et al., 2025; Russinovich and Salem, 2025)
Table 6. Overview of general benchmarks widely used in unlearning evaluation.
4.2.3. Unlearning Robustness
Empirical studies indicate that many machine methods merely suppress the surface-level expression of specific knowledge, leaving the underlying representations vulnerable to various adversarial attacks (Lynch et al., 2024; Hong et al., 2024; Łucki et al., 2025). To systematically assess robustness, a range of adversarial techniques from the security domain have been integrated into the evaluation of LLM unlearning (Lynch et al., 2024; Hong et al., 2024; Li et al., 2024; Che et al., 2025; Sanyal and Mandal, 2025; Jin et al., 2024; Shi et al., 2025; Ramakrishna et al., 2025; Jia et al., 2024), which we collectively refer to as attack techniques. Commonly adopted methods include: (1) Attack on the input, such as crafted jailbreak prompts (Shen et al., 2024a), in-context learning adversarial attacks (Wei et al., 2023), GCG (Zou et al., 2023), AutoPrompt (Shin et al., 2020), BEAST (Sadasivan et al., 2024), PAIR (Chao et al., 2025), persona modulation (Shah et al., 2023), JailbreakHub (Shen et al., 2024b), many-shot jailbreaking (Anil et al., 2025). (2) Attack on the hidden layer, such as probing techniques (Belinkov, 2022; Alain and Bengio, 2016), soft-prompt-based threats (Schwinn et al., 2024; Che et al., 2025), AnonAct (Seyitoğlu et al., 2024), Logit Lens (nostalgebraist, 2020). (3) Membership inference attack (MIA), which is a privacy attack that determines whether specific data samples are part of a model’s training set (Shokri et al., 2017), including LOSS (Yeom et al., 2018), Zlib Entropy (Carlini et al., 2021), Min-K% Prob (Shi et al., 2024) and Min-K%++ Prob (Zhang et al., 2025c). Further heuristic attacks have also been proposed for specific unlearning scenarios (Ashuach et al., 2025; Shen et al., 2025; Du et al., 2025).
In response to the characteristics of LLM unlearning tasks, a relatively unique robustness evaluation method, relearning (Lynch et al., 2024; Joshi et al., 2024), is also frequently used. Relearning evaluates an unlearned model by exposing it to a limited subset of the unlearned data. In in-context relearning, knowledge related to the unlearn set, such as book summaries or relevant background information, are included in the prompts when evaluating the unlearned LLM. When relearning by fine-tuning, the model is full-parameter or LoRA fine-tuned on a small portion of the unlearn set or a related set. Empirical studies consistently demonstrate that relearning can substantially degrade unlearning quality, causing the model to rapidly recapitulate a significant portion of unlearned knowledge from sparse cues (Lynch et al., 2024; Lermen and Rogers-Smith, 2024; Zhan et al., 2024; Guo et al., 2025; Joshi et al., 2024), or begin to systematically avoid generating content related to the unlearning target, even when contextually prompted (Joshi et al., 2024). More critically, Doshi and Stickland (2024) reveal that fine-tuning on an entirely benign dataset can also reverse the unlearning effects, restoring model performance to a level comparable to its state before unlearning.
4.2.4. Unlearning efficiency
While the majority of existing studies concentrate on the efficacy of unlearning, the resource overhead of deploying such algorithms under real-world constraints remains a crucial consideration, including both memory occupation (Bannihatti Kumar et al., 2023; Jia et al., 2024) and computational time consumption. For time cost, a straightforward approach is to directly measure the algorithm’s runtime during experiments (Chen and Yang, 2023; Chowdhury et al., 2025; Bannihatti Kumar et al., 2023; Jia et al., 2024). Since computational speed is closely tied to GPU performance, some studies convert raw runtime into GPU hours (i.e., number of GPUs × training hours) to facilitate comparison (Ren et al., 2025; Ji et al., 2024). Nonetheless, fair cross-study comparisons remain challenging due to variations in experimental environments. The most reliable method is to execute all algorithms under controlled conditions, though this is often resource-intensive. Alternatively, several works estimate time consumption theoretically, using metrics such as floating-point operations (Cha et al., 2024) or gradient computation budgets (Wei et al., 2025b). However, discrepancies between theoretical estimates and actual runtime may arise due to differences in implementation and hardware optimization.
5. Challenges and Future Directions
5.1. Challenges
5.1.1. Definition and Evaluation of Unlearning
In Section 2.1, we characterize the goal of LLM unlearning as ensuring that “the unlearned model should no longer memorize information from the unlearn set while preserving all other knowledge.” However, two key issues remain ambiguous, leading to divergent definitions of unlearning across the literature and, consequently, to inconsistent and imprecise evaluation practices.
(1) How should memorization be defined and detected? Most studies assess memorization based on model output in specific tasks, yet disagree on the criteria for judging these outputs. For the content related to the unlearn set, some works argue that the model should simply avoid generating such content (Eldan and Russinovich, 2024), while others require it to explicitly respond with “I don’t know” (Shen et al., 2025). Another line of research proposes that the unlearned model should produce outputs similar to those of a hypothetical retrained model, such as giving a specific incorrect answer (Rezaei et al., 2024). When direct output is insufficient, adversarial methods are sometimes employed to expose memorization. However, such approaches face inherent limitations: Overly weak attacks may fail to detect memorization, whereas overly strong ones can force the model to generate arbitrary content, casting doubt on their reliability as auditing tools (Chen et al., 2025). Moreover, certain attack methods are considered ill-suited to the LLM context, such as MIA, which typically requires training numerous shadow models, thus being both data-prohibitive and computationally intractable for LLMs (Liu et al., 2024a).
(2) What should constitute the unlearn set? For synthetic datasets such as TOFU (Maini et al., 2024), this question is relatively straightforward. However, in real-world scenarios, data interconnectivity complicate the identification of appropriate unlearning targets. Tian et al. (2024) adopt a legal perspective to determine which copyrighted or private data should be unlearned, while Wu et al. (2025) construct relationship graphs to identify necessary data for removal. Despite their merits, these methods remain reliant on manual, domain-specific analysis, and lack generalizability.
5.1.2. Effect of Unlearning
Unlearning has different effects on different languages and data, further increasing the difficulty of designing and evaluating unlearning algorithms.
Effects across languages. Some studies conduct evaluations with prompts translated into languages other than English, finding that monolingual unlearning is fundamentally insufficient for multilingual LLMs (Lynch et al., 2024; Jin et al., 2024). Furthermore, more languages that systematically divide into high- and low-resource are used in evaluations (Choi et al., 2024a; Lu and Koehn, 2025), revealing the fact that unlearning in one language does not necessarily transfer to others and could even inadvertently reinforce harmful content across languages. Together, these findings underscore a critical consensus: effective and secure unlearning necessitates multilingual joint unlearning strategies that are designed to address model behavior holistically in all languages.
Effects across data. From the perspective of data distribution, Baluta et al. (2024) demonstrate that out-of-distribution (OOD) data require more gradient ascent but offer a better unlearning quality, whereas in-distribution data allow faster unlearning but severely compromise model utility, illustrating a fundamental trade-off between unlearning efficiency and model preservation. Considering the logical connectivity of the data, Choi et al. (2024b) identify that current unlearning methods struggle with multi-hop knowledge, where unlearning one intermediate fact in a chain often fails to remove the entire logical sequence. Furthermore, some studies investigate the impact on adjacent data after performing unlearning on selected data, identifying phenomena called “transfer unlearning” (Lu et al., 2024), “ripple effect” (Zhang et al., 2025d) and “onion effect” (Borkar, 2023). These effects highlight the intricate and unpredictable consequences of unlearning, emphasizing the need for careful monitoring to ensure that unlearning achieves its intended goals without introducing new risks.
5.1.3. Unlearning in Reality
A significant challenge lies in the scaling gap between experimental settings and real-world conditions. Current unlearning experiments are largely limited to models with fewer than 10 billion parameters and unlearning sets under 1 billion instances, raising concerns about the applicability of these methods to larger models and datasets. Shi et al. (2025) analyze how evaluation metrics evolve as the size of the unlearn set increases, providing insight into scalability. On the other hand, in practical deployments, large models are often compressed, such as using quantization for efficiency. Notably, Zhang et al. (2025e) demonstrate that quantizing unlearned models can inadvertently reactivate unlearned knowledge, highlighting a key scalability challenge.
In commercial applications, unlearning requests typically arrive sequentially, requiring models to continuously unlearning while maintaining performance (Rezaei et al., 2024; Veldanda et al., 2024). To assess long-term viability, Shi et al. (2025) collect model checkpoints after processing each sequential request and track evaluation metrics over time. This approach helps quantify the cumulative impact of repeated unlearning and the model’s ability to sustain utility. Unfortunately, current unlearning methods are not yet ready to handle sequential unlearning.
5.2. Future Directions
5.2.1. Unlearning in Specialized Architectures and Scenarios
The field is moving towards addressing unlearning in sophisticated model architectures. Cheng and Amiri (2025) pioneer this effort for tool-augmented large language models (LLMs) by proposing ToolDelete, the first unlearning framework designed to remove a specific “skill” or the ability to use a particular tool, and they introduce a new membership inference attack (MIA) for evaluation. Similarly, the unique structure of Mixture-of-Experts (MoE) models presents a distinct challenge. Zhuang et al. (2025) find that unlearning a single expert is insufficient and propose the Selected Expert Unlearning Framework (SEUF) to effectively perform unlearning on MoE models. These works demonstrate that effective unlearning requires bespoke algorithms tailored to a model’s specific architecture and knowledge organization.
5.2.2. Unlearning as Tools
Unlearning is not only a goal in itself, but also a powerful tool when we further expand the scope of unlearning targets. Firstly, when choosing injected trojans or backdoor triggers as the target, unlearning can be an effective tool in defense (Jiang et al., 2025a; Zhao et al., 2025; Hernandez, 2024). Similarly, an opposite target, such as removing the safety alignment or disrupting the subsequent fine-tuning process on a base model, can convert unlearning to a means of attack (Rashid et al., 2025). Furthermore, if unlearning the selected training data and examining the changes in the model before and after unlearning, we can have novel insights into how different data components contribute to and influence the final model capabilities (Zhao et al., 2024; Isonuma and Titov, 2024). A powerful and accurate unlearning method will play an important role as a tool.
5.2.3. Unlearning beyond Data
Most existing studies focus on unlearning specific data instances. However, in practical scenarios, unlearning requests often target not only concrete data but also abstract concepts or capabilities, such as erroneous reasoning patterns, harmful ethical values, or unsafe skills (Lang et al., 2025; Li et al., 2024). Extending unlearning beyond the data to encompass abstract constructs is essential to prevent the propagation of incorrect or harmful knowledge. Achieving this goal may present two main pathways: one is to precisely identify and modify parameters or representations associated with particular concepts or abilities; the other leverages established alignment techniques, such as reinforcement learning, by designing appropriate reward mechanisms that penalize the generation of undesirable content.
5.2.4. Robust Unlearning
In light of the observed fragility of LLM unlearning, a significant research direction aims to develop techniques that enhance its robustness and long-term stability. These defensive efforts pursue two primary objectives: (1) to ensure that knowledge removal is thorough and persistent, thereby resisting attempts at recovery; and (2) to prevent the unlearning procedure from introducing new vulnerabilities or unintended side effects. Several existing studies address the first objective through robust unlearning frameworks (Tamirisa et al., 2025; Zhang et al., 2024a; Huu-Tien et al., 2025; Fan et al., 2025) or methods that strengthen the robustness of unlearned models (Zhang et al., 2024a; Huu-Tien et al., 2025). Nevertheless, given the proliferation of advanced attacks, achieving truly robust unlearning remains a critical and ongoing topic.
5.2.5. Verifiable and Certifiable Unlearning
In most current practices, unlearning is applied to models that have already internalized the content targeted for removal through opaque mechanisms, complicating the certification of unlearning effectiveness. However, from legal, safety, and social trust perspectives, achieving verifiable and trustworthy unlearning remains critically important. To validate existing unlearning methods, it is essential to establish a fair and comprehensive evaluation benchmark. Looking ahead, future work may also draw inspiration from frameworks such as SISA by designing structured data storage and training protocols to enable intrinsically verifiable unlearning.
6. Conclusions
Machine unlearning has emerged as a pivotal technique to address critical challenges in large language models, including privacy protection, copyright compliance, and safety enhancement. In this survey, we provide a comprehensive review of work dedicated to LLM unlearning, including the definition and goal of LLM unlearning, the most recent LLM unlearning methods, and commonly used datasets and evaluation metrics of unlearning. Despite significant progress, the field of LLM unlearning remains in its early stages, with fundamental challenges in the definition, evaluation, effects and practical deployment of unlearning. Furthermore, we suggest several promising directions for future research. We hope that this survey can provide readers with a general understanding of recent progress in this field and shed some light on future developments.
Acknowledgments
This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0680201), the Beijing Natural Science Foundation (QY23156, 4252023).